OpenStudy (anonymous):

Use the distributive property to simplify the expression below: -4(3u - 5) -12u - 5 -12u + 20 -12u - 20 -12u + 9 Which of the following ordered pairs is a solution to the inequality y > -x + 5? (1, -10) (0, 5) (5, -1) (1, 9)

OpenStudy (anonymous):

answer to the above would be, -4*3*u +5*4 =-12u+20 and to the below, by plugging values is (1,9)

OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6 slope = 5/3; y-intercept = (0, 2) slope = -5; y-intercept = (0, -6) slope = -3/5 ; y-intercept = (0, 2) slope = 5/3; y-intercept = (0, -6)